Grade 11General Physics

a drunker walking in a narrow lane takes 8 steps forward and 6 steps backward and so on.if every step are 1 metre long and required 1 second.How long he wil take fall in a pit 18 metre away from the start?

Bipin Jung Thakuri
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
8 steps(F)- 6 steps(B) = 2 meters forward
2+ 8(F) -6 (B) = 4 meters forward
4+ 8(F) -6 (B) = 6 meters forward
6 + 8 (F) - 6 (B) = 8 meters forward
8 + 8 (F) - 6 (B) = 10 meters forward
10 + 8 (F) = 18 meters; fallen into a pit
Total steps = 78 steps
Hence after 78 seconds he will fall into a pit
 
Regards
Arun (askIITians forum expert)